# Tradeoffs Between Alignment and Helpfulness in Language Models

### 29 Jan 2024 v1

@@ 1 Introduction
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021);  Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019);  Yu & Sagae (2021)Xu et al. (2021)Subhash (2023)Zou et al. (2023b) and the work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, in order to steer from undesired to desired behaviors. With it, they demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023), Leong et al. (2023), Liu et al. (2023b) and Turner et al. (2023) to improve alignment and reduce toxicity.
@@ A.1 Introduction of assumptions
That is to say, that on the axis defined by $\delta r_{e}(q)$, positive and negative representations can be separated, and the minimal distance between representations of positive and negative answers is $\Delta$. It is used in theorm 1, to obtain that the probability of the aligned answers increases w.r.t.the misaligned answers as the coefficients of the injected representations increases.
@@ A.2 Discussion of assumptions
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[\langle v,r_{good}-r_{bad}\rangle]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space.
Normal token logit change (assumption 3): Random vectors in high dimension tend to have a normal distribution around a sphere. If we take an isotropic normal distribution on a sphere and look at the coordinate-wise marginal probability, then it too will look Gaussian. As can be seen in A.3, this assumption is satisfied in real LLMs.
@@ A.3 Experiments for assumptions
Normal assumption
We checked that the change in logit distribution is normal (figure 9) as proposed in assumption 3 and verify that indeed it is approximately normal with the standard deviation growing linearly with the injected vector norm. In figure 9(a) and (b) we show the density of change in logits due to representation engineering on the dataset "international law" from MMLU for the PCA vectors of fairness and harmfulness with different coefficients. In figure 9(c) and (d) we show that the standard deviation of the distributions obtained in (a) and (b) increases linearly with the size of the coefficient. Notice that here we show $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$ distributes normally instead of $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, since if the latter distributes normally, then the former is a subtraction of two normal distributions and should also be normal. The benefit of using the former instead of the latter is that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail, which can lead to an increase in the correct answer’s probability as $|\delta r_{e}(q)|$ is increased.
@@ Appendix B Proof of theorem 1
$=\frac{\sum_{a_{+}\in good}exp(<r(q)+\delta r(q),U^{T}e_{a_{+}}>)-\sum_{a_{-}\in bad}exp(<r(q)+\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q)+\delta r(q),U^{T}e_{a_{+}}>)+\sum_{a_{-}\in bad}exp(<r(q)+\delta r(q),U^{T}e_{a_{-}}>)}=$ (15)
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers.
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(<r(q)+\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q)+\delta r(q),U^{T}e_{a_{+}}>)}}{1+\frac{\sum_{a_{-}\in bad}exp(<r(q)+\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q)+\delta r(q),U^{T}e_{a_{+}}>)}}=$ (16)
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)exp(<\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}>)exp(<\delta r(q),U^{T}e_{a_{+}}>)}}{1+\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)exp(<\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}>)exp(<\delta r(q),U^{T}e_{a_{+}}>)}}=$ (17)
@@
$\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)exp(<\delta r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}>)exp(<\delta r(q),U^{T}e_{a_{+}}>)}<$ (18)
$<\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)\cdot min_{a^{\prime}_{-}\in bad}\{exp(<\delta r(q),U^{T}e_{a^{\prime}_{-}}>)\}}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}>)\cdot max_{a^{\prime}_{+}\in good}exp(<\delta r(q),U^{T}e_{a^{\prime}_{+}}>)}=$ (19)
$=\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}>)}\cdot\frac{1}{max_{a^{\prime}_{+}\in good,a_{-}\in bad}exp(<\delta r(q),U^{T}e_{a^{\prime}_{+}}>)}$ (20)
$=\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}>)}\cdot\frac{1}{exp(max_{a^{\prime}_{+}\in good,a_{-}\in bad}<\frac{\delta r(q)}{|\delta r(q)|},U^{T}e_{a^{\prime}_{+}}>\cdot|\delta r(q)|)}$ (21)
From $\Delta$ margin SVM classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$:
$<\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}>)}\cdot\frac{1}{exp(\Delta|\delta r|)}$ (22)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}>)}\cdot\frac{1}{exp(\Delta|\delta r|)}}{1+\frac{\sum_{a_{-}\in bad}exp(<r(q),U^{T}e_{a_{-}}>)}{\sum_{a_{+}\in good}exp(<r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}>)}\cdot\frac{1}{exp(\Delta|\delta r|)}}=$ (23)
$=\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in bad}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in bad}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}$ (24)
$=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in bad}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (25)
Lastly, notice that:
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in bad}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (26)
@@
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$, hence:
@@ Appendix C Proof of theorem 2
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$, hence:
@@
Under the assumption that $X_{i}$ distribute normally, $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions. Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.

### 05 Feb 2024 v2

@@ 1 Introduction
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021); Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b) and the work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, in order to steer from undesired to desired behaviors. With it, they demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023b); Turner et al. (2023) to improve alignment and reduce toxicity.
@@ 4 Main Results
Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in E, to yield similar results.
@@ 6 Discussion
Overall, we hope that our theoretical work will shed light on the mechanism of representation engineering, which provides a new interesting direction for language model alignment.
@@ A.1 Introduction of assumptions
That is to say, that on the axis defined by $\delta r_{e}(q)$, positive and negative representations can be separated, and the minimal distance between representations of positive and negative answers on it is $\Delta$. It is used in theorm 1, to obtain that the probability of the aligned answers increases w.r.t.the misaligned answers as the coefficients of the injected representations increases.
Note that the above assumption can be relaxed from a hard margin to a soft margin assumption, where $\delta r_{e}(q)$ classifies the representations of positive and negative answers, but part of the misaligned/aligned answers’ representations are misclassified as aligned/misaligned. This yields similar results to theorem 1 that are shown in appendix E.
@@ A.2 Discussion of assumptions
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[\langle v,r_{good}-r_{bad}\rangle]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix E we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Normal token logit change (assumption 3): Random vectors in high dimension tend to have a normal distribution around a sphere. If we take an isotropic normal distribution on a sphere and look at the coordinate-wise marginal probability, then it too will look Gaussian. The change to logits does not need to be exactly normal for theorem 2 to hold, it suffices that on average the correct answer logit does not increase relative to the incorrect answers’ logits (i.e.  some incorrect answer logits are enhanced w.r.t.the correct answer logit due to representation engineering and some are reduced), but we choose this assumption due to the theoretical motivation and as can be seen in A.3, this assumption is approximately satisfied in real LLMs.
@@ A.3 Experiments for assumptions
Normal assumption
We checked that the change in logit distribution is normal (figure 10) as proposed in assumption 3 and verify that indeed it is approximately normal with the standard deviation growing linearly with the injected vector norm. Note that the latter condition is because in assumption 3, we normalize $\delta r_{e}(q)$ before taking the inner product in equation 12, while here we do not, thus we expect to see the standard deviation of the Gaussians grow linearly with $|\delta r_{e}|$ which as shown above grows linearly with $r_{e}$. In figure 10(a) and (b) we show the density of change in the top ten token’s logits due to representation engineering on the dataset "international law" from MMLU for the PCA vectors of fairness and harmfulness with different coefficients. In figure 10(c) and (d) we show that the standard deviation of the distributions obtained in (a) and (b) increases linearly with the size of the coefficient. Notice that here we show $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$ distributes normally instead of $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, since if the latter distributes normally, then the former is a subtraction of two normal distributions and should also be normal. The benefit of using the former instead of the latter is that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail, which can lead to an increase in the correct answer’s probability as $|\delta r_{e}(q)|$ is increased.
@@ Appendix B Proof of theorem 1
$=\frac{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)-\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)+\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}=$ (15)
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively.
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r(q),U^{T}e_{a_{+}}\rangle)}}=$ (16)
$=\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}}=$ (17)
@@
$\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)exp(\langle\delta r(q),U^{T}e_{a_{+}}\rangle)}<$ (18)
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)\cdot\min_{a^{\prime}_{-}\in bad}\{exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{-}}\rangle)\}}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)\cdot\max_{a^{\prime}_{+}\in good}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}=$ (19)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{\max_{a^{\prime}_{+}\in good,a_{-}\in bad}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}$ (20)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\max_{a^{\prime}_{+}\in good,a_{-}\in bad}\langle\frac{\delta r(q)}{|\delta r(q)|},U^{T}e_{a^{\prime}_{+}}\rangle\cdot|\delta r(q)|)}$ (21)
From $\Delta$ margin SVM classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 2):
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}$ (22)
@@
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}{1+\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\Delta|\delta r|)}}=$ (23)
$=\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)}$ (24)
$=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (25)
Then, notice that:
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (26)
@@
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
@@ Appendix C Proof of theorem 2
Lastly, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
@@
Under the assumption that $X_{i}$ distribute normally (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions. Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
@@ D.3 Helpfulness experiments
Appendix E Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to representation engineering SVM classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
$i\in aligned,j\in misaligned:~{}~{}~{}\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\leq\Delta$ (38)
That is, the margin $\Delta$ does not hold for every pair of aligned and misaligned answers.
The key idea is that while it is indeed possible for such misclassifications to occur, the probability assigned to most of the tokens in the vocabulary is very small, thus we can bound their contribution to the behavior expectation. To this end, we define a set of misclassified responses: $\{i\in misclassified\}$ and bound the probability mass that the model assigns them by:
$\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ (39)
Furthermore, we bound how "deep" the misclassified negative response representations can go into the cluster of positive answer representations:
$\min_{i\in aligned,j\in misclassified}\{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\}>-M$ (40)
With this, the SVM classification assumption can be modified as:
Assumption 4.
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , SVM classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma^{\star}\rightarrow\{-1,+1\}$:
$\min_{i\in aligned,j\in misaligned}\bigg{\{}\big{\langle}\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}-U^{T}e_{j}\big{\rangle}\bigg{\}}>\Delta$ (41)
Up to a set of misclassified answers, whose probability is bounded by $\sum_{i\in misclassified}\mathbb{P}_{\theta}(i|q)<\delta\cdot\sum_{i\in aligned}\mathbb{P}_{\theta}(i|q)$ that satisfy:
$\min_{i\in aligned,j\in misclassified}\{\langle\delta r_{e}(q),U^{T}(e_{i}-e_{j})\rangle\}>-M$ (42)
Note that realistically, $\delta$ can be very small for a very large set of tokens, as in inference, LLMs typically assign high probability to few tokens and very low probability for most. Hence it suffices to classify just a few high probability tokens.
We can restate theorem 1 in the following way:
Theorem 3.
Let $\delta,\epsilon>0$ and let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{\star}\rightarrow\{-1,+1\}$ be a behavior scoring function. Under assumption 4, for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$ the behavior expectation of the model conditioned on the query $q$ satisfies:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]\geq tanh(\Delta\lambda\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (43)
Proof.
We follow the proof of theorem 1, up to equation 24, there, we introduce the misclassified tokens’ contributions, which we denote by $R=\frac{\sum_{a\in misclassified}exp(\langle r(q)+\delta r_{e}(q),U^{T}e_{a})}{\sum_{a_{+}\in good}exp(\langle r(q)+\delta r_{e}(q),U^{T}e_{a_{+}})}$:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{1-\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)-R}{1+\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}exp(-\Delta|\delta r|)+R}$ (44)
Following the same idea as with equation 18, we obtain that:
$R<\frac{\sum_{a\in misclassified}exp(\langle r(q),U^{T}e_{a})}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}})}\frac{1}{exp(-|\delta r|M)}$ (45)
Plugging this in gives:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>\frac{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)-\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)+\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}>$ (46)
Denote the first second and third terms respectively as $A,B,C$:
$=\frac{A-B-C}{A+B+C}=\frac{\frac{A-B}{A+B}-\frac{C}{A+B}}{1+\frac{C}{A+B}}>(\frac{A-B}{A+B}-\frac{C}{A+B})(1-\frac{C}{A+B})>\frac{A-B}{A+B}-2\frac{C}{A+B}$ (47)
Notice that from the transition in equation 25:
$\frac{A-B}{A+B}=tanh(\frac{\Delta|\delta r|-\ln(\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)})}{2})$ (48)
Is the bound from theorem 1, and the second term:
$\frac{C}{A+B}=\frac{\sum_{a\in misclassified}\mathbb{P}_{\theta}(a|q)exp(M|\delta r|)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)+\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)exp(-\Delta|\delta r|)}<\delta\cdot exp(M|\delta r|)$ (49)
Lastly, notice that:
$\frac{\sum_{a_{-}\in bad}\mathbb{P}_{\theta}(a_{-}|q)}{\sum_{a_{+}\in good}\mathbb{P}_{\theta}(a_{+}|q)}=\frac{1-B_{0}}{1+B_{0}}$ (50)
Where $B_{0}=B[\mathbb{P}_{\theta}(\cdot|q)]$, and that:
$arctanh(B_{0})=-\frac{1}{2}\ln\frac{1-B_{0}}{1+B_{0}}$ (51)
Thus we obtain:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta|\delta r(q)|}{2}+arctanh(B_{0}))-2\delta\cdot exp(M|\delta r|)$ (52)
Then, note that for coefficients that are not too large, $|\delta r(q)|$ is proportional to the injected vector coefficient $r_{e}$ (assumption 1), hence:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-2\delta\cdot exp(M\lambda\cdot r_{e})$ (53)
Where $\lambda$ is the coefficient relating $r_{e}$ to $|\delta r(q)|$. Thus for $r_{e}<\frac{\log\frac{\epsilon}{2\delta}}{M\cdot\lambda}$:
$B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>tanh(\frac{\Delta\lambda}{2}\cdot r_{e}+arctanh(B_{0}))-\epsilon$ (54)